OpenStudy (anonymous):

what is the simplified form of b/2 divided by 4/2b?

OpenStudy (anonymous):

\[\frac{ \frac{ b }{ 2 } }{ \frac{ 4 }{ 2b } }\]

OpenStudy (anonymous):

the 2b flips up and becomes the numerator, 4 stays as the denominator

OpenStudy (anonymous):

\[\frac{ b }{ 2 }\frac{ 2b }{ 4 }\]

OpenStudy (anonymous):

can you simplify this?

OpenStudy (anonymous):

Is it cross multiplying?

OpenStudy (anonymous):

not quite; multiply the numerators and denominators together

OpenStudy (anonymous):

so what do you get?

OpenStudy (anonymous):

2b^2/8

OpenStudy (anonymous):

\[\frac{ 2b ^{3} }{ 8 }\]

OpenStudy (anonymous):

then wat?

OpenStudy (anonymous):

my bad it should be b^2 instead of b^3

OpenStudy (anonymous):

that's your ans

OpenStudy (anonymous):

are u sure?

OpenStudy (anonymous):

wait, so it's b^2/4 right?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

YAY thanks